OpenStudy (anonymous):

how do you solve 3/2x-2 + 1/2 = 2/x-1

OpenStudy (anonymous):

\[\frac{3}{2x}-2+\frac{1}{2} = \frac{2}{x}-1\] Is this the problem?

OpenStudy (anonymous):

the minus 2 & 1 go with the 2x & x

OpenStudy (anonymous):

\[\frac{3}{2x-2}+\frac{1}{2} = \frac{2}{x-1}\] Right click on the equation and go to "Show Source" to see how this is typed

OpenStudy (anonymous):

okaaaay . how do you do it now

OpenStudy (anonymous):

gotcha, first things first, multiply everything by 2 to make it a little nicer to look at this should give you: 3/(x-1) +2 = 4/(x-1)

OpenStudy (anonymous):

from here you can notice that we can write the 3/(x-1) and the 4/(x-1) as 3(1/(x-1)) and 4(1/(x-1))

OpenStudy (anonymous):

since these are like terms you can add them together.

OpenStudy (anonymous):

subtracting 3(1/(x-1)) from each side, you should get: 2=1/(x-1)

OpenStudy (anonymous):

now multiply each side by (x-1) to take care of the denominator and get: 2x-2 = 1

OpenStudy (anonymous):

add the 2 to each side: 2x = 3 and divide by 2 x= 3/2

OpenStudy (anonymous):

Combine the two fractions on the left side by multiplying the top and bottom of the (1/2) by (x-1) so they have a common denominator:\[\frac{3+(x-1)}{2x-2} = \frac{2}{x-1}\]Then cross multiply and arrange the equation so it is equal to zero because it will be a polynomial:\[(x-1)(3+x-1) = 2(2x-2)\]\[x^{2}-3x+2=0\]Then factor and solve for zero:\[(x-2)(x-1) = 0\]So x must equal 2 or 1.